3. A circular steel wire 2.00 m long must stretch no more than 0.25 cm when a tensile force of 400 N is applied to each end of the wire. What minimum diameter is required for the wire?

1
Expert's answer
2022-06-07T17:37:26-0400

"d=\sqrt\frac{4Fl}{\pi E∆l}=0.2~mm."